Use the Laplace transform to solve the given initial-value problem.

y' − y = 2 cos(4t),  y(0) = 0

y(t) =


1
Expert's answer
2021-04-13T13:29:10-0400

Let us use the Laplace transform to solve the given initial-value problem:

"y' − y = 2 \cos(4t),\ y(0) = 0".


Taking into account that "y(t)\to Y(p),\ y'(t)\to pY(p)-y(0), \cos 4t\to\frac{p}{p^2+16}", we have the equation:


"pY(p)-Y(p)=\frac{2p}{p^2+16}" or "Y(p)(p-1)=\frac{2p}{p^2+16}" or "Y(p)=\frac{2p}{(p-1)(p^2+16)}".


It follows that "\frac{2p}{(p-1)(p^2+16)}=\frac{A}{p-1}+\frac{Bp+C}{p^2+16}" and hence


"2p=Ap^2+16A+Bp^2+Cp-Bp-C"


Therefore, we have the following system:


"\begin{cases}
A+B=0\\
C-B=2\\
16A-C=0
\end{cases}"


"\begin{cases}
A+C=2\\
B=C-2\\
16A-C=0
\end{cases}"


"\begin{cases}
17A=2\\
B=C-2\\
C=16A
\end{cases}"


"\begin{cases}
A=\frac{2}{17}\\
B=C-2\\
C=\frac{32}{17}
\end{cases}"


"\begin{cases}
A=\frac{2}{17}\\
B=-\frac{2}{17}\\
C=\frac{32}{17}
\end{cases}"


We conclude that "Y(p)=\frac{2}{17}\frac{1}{p-1}-\frac{2}{17}\frac{p}{p^2+16}+\frac{32}{17}\frac{1}{p^2+16}", and using "\frac{1}{p-1}\to e^t,\ \frac{4}{p^2+16}\to\sin 4t" we have the solution:


"y(t)=\frac{2}{17}e^t-\frac{2}{17}\cos 4t+\frac{8}{17}\sin 4t."